OpenStudy (mendicant_bias):

I'm trying to find the area of the portion of a tilted plane inside a cylinder using parametrization, posted below momentarily.

OpenStudy (mendicant_bias):

http://i.imgur.com/7nHAfPy.png How exactly should I set up the vector r? I'm guessing I should base the parametrization of r off of the cylinder?

OpenStudy (mendicant_bias):

\[R(u,v) = \ ?\]

OpenStudy (mendicant_bias):

@FibonacciChick666 , could you help me out on this? Just trying to figure out the parametrization.

OpenStudy (fibonaccichick666):

I'm working on a paper, but here is a really good explanation someone else wrote http://math.stackexchange.com/questions/817869/surface-integral-the-area-of-a-plane-inside-a-cylinder

OpenStudy (mendicant_bias):

ty!

OpenStudy (fibonaccichick666):

np, good luck on the assignment

OpenStudy (mendicant_bias):

I'm trying to take a shot at this using a calculus-based approach, and have come up with this. @Zarkon , could you help me out with this/does this setup look okay for the parametrization of the curve intersecting the cylinder and the tilted plane? http://i.imgur.com/xhsfoyR.png

OpenStudy (mendicant_bias):

\[R(r,\theta)=r \cos(\theta)i + r \sin(\theta)j + (1-\frac{1}{2}r \sin(\theta))k.\]

OpenStudy (mendicant_bias):

\[R_r=\cos(\theta)+\sin(\theta)j-\frac{1}{2}\sin(\theta)k.\]

OpenStudy (mendicant_bias):

\[R_\theta = -r \sin(\theta)i+r \cos(\theta)j-\frac{1}{2}r \cos(\theta)k\]

OpenStudy (mendicant_bias):

@wio , could you help me out with this when you get a chance? Thanks so far.

OpenStudy (anonymous):

This is a tricky one, no doubt.

OpenStudy (anonymous):

If you use cylindrical coords then...

OpenStudy (mendicant_bias):

Sorry, was working on another problem. Wait, why isn't it 2z = 2 - y?

OpenStudy (anonymous):

\[ y+2z=2\implies z=1-\frac y2 = 1-\frac{r\sin\theta}{2} \]

OpenStudy (mendicant_bias):

\[y+2z=2; \ \ \ 2z = 2-y; \ \ \ z = \frac{2-y}{2}\]

OpenStudy (anonymous):

Either way is fine.

OpenStudy (anonymous):

Your partial derivatives are fine as well.

OpenStudy (mendicant_bias):

Alright, cool, taking a shot at computing the cross product now between the two vectors that come from it. How do I set up the, uh....the determinant matrix in laTeX in this kinda environment? I think I'll just write it out

OpenStudy (anonymous):

\[ \left|\mathbf \Phi_{r}\times \mathbf \Phi_{\theta}\right|= \begin{vmatrix} \mathbf i &\mathbf j &\mathbf k \\ \cos(\theta)&\sin(\theta)&-\frac 12 \sin(\theta)\\ -r\sin(\theta)&r\cos(\theta)&-\frac 12r \cos(\theta) \end{vmatrix} \]

OpenStudy (mendicant_bias):

Whoop, here we go: http://i.imgur.com/rLXc9Mk.png

OpenStudy (mendicant_bias):

(Just double checking the vector before taking the magnitude o fit)

OpenStudy (mendicant_bias):

Alright, now, taking the magnitude looks super ugly....takign a shot at it now.

OpenStudy (anonymous):

Now that I think about it, I wonder if the normal vectors on the plane would be different.

OpenStudy (mendicant_bias):

Hm?

OpenStudy (anonymous):

The normal vectors on the plane should be the same direction as the normal vectors on the disk we parametrized.

OpenStudy (anonymous):

It's just maybe the magnitude is different.

OpenStudy (mendicant_bias):

I dunno, I'm being clueless atm and don't catch your drift; should I be taking the magnitude of that cross product as it stands?

OpenStudy (anonymous):

Just keep doing what you were doing.

OpenStudy (mendicant_bias):

http://i.imgur.com/IVXMubo.png

OpenStudy (mendicant_bias):

Not putting that all under a square root and summing it,

OpenStudy (anonymous):

Okay, how the hell are you getting that?

OpenStudy (mendicant_bias):

Lul

OpenStudy (mendicant_bias):

Do we agree that the vector, before taking its magnitude, is http://i.imgur.com/BTgWUiw.png

OpenStudy (anonymous):

\[ \begin{split} \left|\mathbf \Phi_{r}\times \mathbf \Phi_{\theta}\right| &=& \begin{vmatrix} \mathbf i &\mathbf j &\mathbf k \\ \cos(\theta)&\sin(\theta)&-\frac 12 \sin(\theta)\\ -r\sin(\theta)&r\cos(\theta)&-\frac 12r \cos(\theta) \end{vmatrix} \\ &=&+ \mathbf i\left(-\frac 12\sin(\theta) r\cos(\theta)+\frac 12\sin(\theta) r\cos(\theta)\right)\\ &&-\mathbf j\left(-\frac 12r\cos^2(\theta)-\frac 12r\sin^2(\theta)\right)\\ &&+\mathbf k\left(r\cos^2(\theta)+r\sin^2(\theta)\right) \\ &=&\left|\frac 12 r\mathbf j+r\mathbf k\right| = \sqrt{\frac 14r^2+r^2} = \frac{\sqrt 5}{2}r \end{split} \]

OpenStudy (mendicant_bias):

J component is one minus the other though, right? Minus a minus?

OpenStudy (mendicant_bias):

(but yeah I get your point lol)

OpenStudy (anonymous):

The term is positive, but it gets subtracted.

OpenStudy (mendicant_bias):

Wait, the term is *negative*, but it gets subtracted, right, so it makes it a positive. One sec.

OpenStudy (mendicant_bias):

OH, I missed a minus sign in my original rprlbem neverMKFEWOHOJA

OpenStudy (mendicant_bias):

I gotchu

OpenStudy (mendicant_bias):

I'm bad at maths without paper, lol.

OpenStudy (mendicant_bias):

Alright, well, cool, got the general setup. Now going to go work some other problems, just need to watch for algebra errors. Thank you.